OpenStudy (anonymous):

What is the factor of 4-(x+y)^2

OpenStudy (anonymous):

You can break it down to (2-(x+y))(2+(x+y)) = 0. So your factors are x+y=2 and x+y=-2

OpenStudy (anonymous):

Oh ok because the answer that they gave me is 2+x+y

OpenStudy (anonymous):

It has two factors. x+y+2 and x+y-2

OpenStudy (anonymous):

Well it gives me both of those choices but it only says that 2+x+y is correct

OpenStudy (anonymous):

Not sure why. Both of those are factors.